Can someone help me on this problem?
Rose's garden is in the shape of a trapezoid. If the height of the trapezoid is 16 m, one base is 20 m, and the area is 224 m^2, find the length of the other base.
Thanks,
Ashley

Answer by checkley71(8403)

224=16*(20+X)/2 OR 224=8(20+X) OR 244=160+8X OR 8X=244-160 OR 8X=64 OR X=8
THEREFORE THE MISSING SIDE IS 20+8=28
PROOF (20+8)/2*16=224 OR 14*16=224